how do i solve this problem 5a+4(3a-8)=4+13a
Answer by stanbon(75887)

how do i solve this problem 5a+4(3a-8)=4+13a
--------------
5a+12a-32 = 4+13a
4a = 36
a = 9
=======
Cheers,
Stan H.
=========